### v2

# Dissociation-driven quarkonium spin alignment in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV

Abstract
The observation of spin alignment of quarkonia in ultra-relativistic heavy-ion collisions provides deep insight into the possible formation of the quark-gluon plasma (QGP). The present study investigates the spin alignment of quarkonia induced by dissociation mechanisms arising from medium effects imposed on quarkonia. We implement an effective Hamiltonian with a medium-modified color-singlet potential to incorporate the coupling of quarkonium spin with medium vorticity. This coupling gives rise to spin-dependent dissociation, which we identify as a plausible mechanism contributing to quarkonium spin alignment. Within the ambit of second-order relativistic viscous hydrodynamics, we calculate the spin-dependent decay widths of charmonium ($J/\psi$, $\psi$(2S)) and bottomonium ($\Upsilon$(1S), $\Upsilon$(2S)) in a rotating thermal medium, including collisional damping and gluonic dissociation effects. We evaluate the observable $\rho_{00}$ for Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV as a function of transverse momentum of the quarkonia, charged particle multiplicity, and medium rotation. The results demonstrate that medium vorticity modifies the quarkonia net decay width and, as a consequence, quarkonia spin alignment gets modified. These findings suggest new directions for understanding spin transport and the microscopic dynamics of vortical QGP.
I Introduction
The hunt for the deconfined phase of strongly interacting matter known as quark-gluon plasma (QGP) in ultra-relativistic heavy-ion collisions was at its peak, but then something unexpected happened in the grand caves under the Jura mountains. The evidence for QGP formation in heavy-ion collisions has been sought through observables such as collective flow, strangeness enhancement, quarkonia suppression, etc. [1, 2, 3, 4, 5]. Similar to those, a substantial influence of the medium on observables, like strangeness enhancement and collective flow at high multiplicities in ppp-p collisions, was observed at the LHC energies [6, 7, 8]. These findings suggest the possible emergence of QGP-like phenomena in small systems, thereby opening a new research domain to investigate the existence of QGP in such small collision systems. These findings also pose a challenge to established methodologies, particularly for observables that use ppp-p collisions as a baseline, like quarkonia suppression. However, quarkonia, being the bound states of the heavy quark and antiquark pair, have masses significantly larger than the QCD scale, $\Lambda_{\rm QCD}$. The large heavy quark masses originate mainly from their Yukawa couplings to the Higgs field through electroweak symmetry breaking, in contrast to light quarks whose effective masses are largely generated by non-perturbative QCD dynamics and spontaneous chiral symmetry breaking. Therefore, quarkonia, with their unique properties, such as large binding energies, small spatial radii, and distinct sequential melting temperatures, make them a better probe for investigation of deconfined QCD matter  [1]. Since quarkonium states are produced predominantly during the early hard scatterings, the vorticity produced in the earliest times of collisions is expected to affect heavy quarks (antiquarks) more strongly than light flavors. Through spin-orbit coupling, such early-time vorticity can induce polarization in quarkonium states, which would access the full space–time evolution of the system, and may provide a direct connection between the early dynamics and the final states.
Recent experimental measurements of hyperon spin polarization and vector meson spin alignment have initiated a more focused study on hadron polarization in ultra-relativistic heavy-ion collisions [9, 10, 13, 11, 12, 14]. Substantial progress has been made to comprehend global, local, and transverse spin polarization of various hadrons from both theoretical and experimental standpoint [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. In this context, the first measurement of $J/\psi$ spin alignment in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV has revealed considerable polarization signals in multiple reference frames, including the helicity, Collins-Soper, and event-plane frames  [28, 29]. These observations have stimulated considerable theoretical interest, motivating efforts to make predictions consistent with experimental results and to identify the possible underlying mechanisms of $J/\psi$ polarization. Among the proposed sources, the vorticity field generated in peripheral collisions is of particular importance. Arising from the large initial orbital angular momentum (OAM) of the colliding nuclei, this vorticity can polarize quarks through spin-orbit coupling, with the resulting polarization subsequently transferred to the produced hadrons  [15, 16].
The spin alignment of quarkonia is characterized through the elements of a $3\times 3$ Hermitian spin density matrix, $\rho_{m,m^{\prime}}$, where $m$ and $m^{\prime}$ denote the spin projections along a chosen quantization axis. The diagonal elements $\rho_{11}$, $\rho_{00}$, and $\rho_{-1-1}$ represent the probabilities of finding the vector meson in spin states +1+1, 0, and 1-1, respectively. Among these, $\rho_{00}$ is of particular importance, as it can be directly measured in the experiments through the angular distribution of the decay products. For quarkonia, $\rho_{00}$ is determined from the angular distribution of decay leptons in the quarkonium rest frame, which is given  [25];
$\frac{dN}{d\Omega}\propto 1+\lambda_{\theta}\cos^{2}\theta+\lambda_{\phi}\sin^{2}\theta\cos 2\phi+\lambda_{\theta\phi}\sin 2\theta\cos\phi$ (1)
where $\theta$ and $\phi$ are the polar and azimuthal angles of the decay products, while $\lambda_{\theta}$, $\lambda_{\phi}$ and $\lambda_{\theta\phi}$ are the polarization parameters that quantify the anisotropy of the decay distribution. These parameters are directly related to the elements of the spin density matrix. The parameter $\rho_{00}$, which quantifies the probability of the vector meson being in a spin-zero projection along the quantization axis, is connected to the polar anisotropy parameter $\lambda_{\theta}$ by the relation,
$\rho_{00}=\frac{1-\lambda_{\theta}}{3+\lambda_{\theta}}$ (2)
In the absence of any spin alignment, the three spin substates ($m_{j}=+1,0,-1$) are equally probable, giving $\rho_{00}=1/3$. The deviation $\rho_{00}-1/3$ thus quantifies the degree of alignment of the vector meson. A value $\rho_{00}<1/3$ indicates a transverse alignment, whereas $\rho_{00}>1/3$ corresponds to a longitudinal alignment with respect to the chosen quantization axis. A precise determination of $\rho_{00}$ and its associated polarization parameters in different reference frames, such as the helicity, Collins–Soper, and event-plane frames, provides key insights into the spin dynamics of quarkonia and the influence of the surrounding medium in relativistic heavy-ion collisions. However, the recent experimental measurement of spin alignment of $J/\psi$ in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV is not consistent across different polarization frame of reference  [28, 29]. The spin alignment observable $\rho_{00}<$ 1/3 for helicity and event plane frame, while it is observed to be $\rho_{00}>$ 1/3 Collins-Soper frame.
In this study, we investigate the spin alignment of charmonium and bottomonium states, specifically $J/\psi$, $\psi$(2S), $\Upsilon$(1S), and $\Upsilon$(2S) as a function of transverse momentum ($p_{\rm T}$) and charged particle multiplicity in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02~\text{TeV}$. This study examines the spin alignment of quarkonia along the vorticity or total angular momentum direction, which is perpendicular to the reaction plane [30]. In the absence of fluctuations, the event plane serves as an estimate of the reaction plane [31]. Therefore, the present calculation is more closely aligned with the event-plane frame. Our study focuses on the spin-dependent dissociation of quarkonia as a possible mechanism for their spin alignment in a vortical QGP medium. Here, we have obtained the spin alignment of quarkonium states considering collisional damping and the gluonic dissociation mechanisms. The combined effect of these in-medium mechanisms on the spin alignment observable $\rho_{00}$ is systematically explored. Further, the space-time evolution of the QGP is modeled using a second-order relativistic viscous hydrodynamic framework, which provides the temperature cooling profile in the expanding medium. The effect of medium vorticity, characterized through the conserved circulation $C$, is further examined to assess its role in generating spin alignment for different quarkonium states within the deconfined phase of thermally equilibrated hot and vortical QCD matter.
The paper is organized as follows: Section II presents the detailed theoretical formulation used in this study. Section III discusses the results and their physical implications. Finally, Section IV summarizes our findings and provides an outlook for future research.
II Formulation
The evolution of quarkonium states within the QGP medium is described using the time-independent Schrödinger equation (SE) and followed by dissociation mechanisms like collisional damping and gluonic dissociation. Collisional damping and gluon-induced dissociation are among the extensively studied processes governing quarkonium suppression in the QGP medium [35, 36, 32, 37, 33, 34]. These mechanisms induce spin-dependent modifications in their dissociation widths, thereby influencing the spin density matrix element, $\rho_{00}$. To quantify this effect, we calculate the spin-dependent decay widths arising from both collisional damping and gluonic dissociation by incorporating the coupling between the spin of the particle and the vorticity of the medium within an effective Hamiltonian framework that includes a medium-modified potential. Given the pivotal role of temperature in modifying the quarkonia characteristics within the QGP medium, it is imperative first to examine the thermodynamic evolution of deconfined QCD matter.
II.1 Thermal Evolution of the Medium
The possible formation of QGP in ultra-relativistic heavy-ion collisions undergoes rapid expansion in both space and time domains due to strong internal pressure gradients. The space–time evolution of such a deconfined medium is proposed to be governed by relativistic hydrodynamics. In the present study, we employ the second-order relativistic viscous hydrodynamics framework to describe the dynamical evolution of the QGP. To solve the hydrodynamic equations, one must specify appropriate initial conditions and an equation of state (EoS) that connects with the thermodynamic quantities of the system. In the absence of a first-principles determination of the initial temperature ($T_{0}$) and thermalization time ($\tau_{0}$) of QGP, we estimate $T_{0}$ using a phenomenological relation [Hwa:1984fj] that connects it to the experimentally measured charged particle multiplicity:
$T_{0}=\left[\frac{90}{g_{k}4\pi^{2}}C^{\prime}\frac{1}{A_{T}\tau_{0}}\frac{dN_{\text{ch}}}{dy}\right]^{1/3},$ (3)
here the constant is $C^{\prime}=\frac{2\pi^{4}}{45\zeta(3)}\approx 3.6$, and $\frac{dN_{\text{ch}}}{dy}\simeq\frac{dN_{\text{ch}}}{d\eta}$, under the massless limit. Equation 3 is derived considering the isentropic expansion of the fireball. In a QGP medium with comparatively small shear viscosity, the uncertainties introduced by this assumption are negligible compared to those arising from other parameters such as the transverse overlap area ($A_{T}$), thermalization time, and statistical degeneracy ($g_{k}$) of the deconfined phase. The transverse overlap area, $A_{T}=\pi R_{T}^{2}$, is estimated using the MC-Glauber mode for Pb–Pb collision at $\sqrt{s_{\rm NN}}=5.02$ TeV, where $R_{T}$ is the transverse radius of the fireball. It is important to emphasize that the initial thermalization time $\tau_{0}$ plays a decisive role in determining the initial temperature $T_{0}$. In the present work, we adopt $\tau_{0}=0.2~\text{fm}/c$ for Pb–Pb collisions at $\sqrt{s_{\text{NN}}}=5.02~\text{TeV}$, the chosen value is consistent with those used in modern viscous hydrodynamic simulations. Corresponding to this choice of $\tau_{0}$, the values of $T_{0}$ are obtained as a function of the measured charged particle multiplicity. The charged particle multiplicity classes for Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV is taken from ALICE experimental data [39] and reported in Table 1.
Multiplicity Class Centrality Class $\langle\frac{dN_{\text{\rm ch}}}{d\eta}\rangle$
Mult. Class - 1 (70-80)% 44.9 $\pm$ 3.4
Mult. Class - 2 (60-70)% 96.3 $\pm$ 5.8
Mult. Class - 3 (50-60)% 183 $\pm$ 8
Mult. Class - 4 (40-50)% 318 $\pm$ 12
Mult. Class - 5 (30-40)% 512 $\pm$ 15
Mult. Class - 6 (20-30)% 786 $\pm$ 20
Mult. Class - 7 (10-20)% 1180 $\pm$ 31
Mult. Class - 8 (7.5-10)% 1505 $\pm$ 44
Mult. Class - 9 (5.0-7.5)% 1666 $\pm$ 48
Mult. Class - 10 (2.5-5.0)% 1850 $\pm$ 55
Mult. Class - 11 (0-2.5)% 2035 $\pm$ 52
Table 1: The $\langle\frac{dN_{\text{\rm ch}}}{d\eta}\rangle$ values measured for Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV in $|\eta|<$ 0.5 for 11 centrality classes are taken from ALICE experimental data [39].
Now, with the initial conditions in hand, we employ 1+1D with boost invariance second-order viscous hydrodynamics to determine the temperature cooling profile in the medium [40]. The second-order cooling law is derived from kinetic theory using Grad’s 14-moment approximation method within the framework of the Müller-Israel-Stewart equation [41, 42, 43, 44], given as;
$\frac{dT}{d\tau}=-\frac{T}{3\tau}+\frac{T^{-3}\phi}{12a\tau}$ (4)
and
$\frac{d\phi}{d\tau}=-\frac{2aT\phi}{3b}-\frac{1}{2}\phi\left[\frac{1}{\tau}-\frac{5}{T}\frac{dT}{d\tau}\right]+\frac{8aT^{4}}{9\tau}$ (5)
The variable $\phi$ in Eq. (5) is called the shear viscous term, which quantifies the time-dependent change in the shear viscosity ($\eta$). In essence, $\phi$ reflects the characteristics of the medium created under extreme conditions during ultra-relativistic collisions. For the first-order solution of Eq. (4), $\phi$ is defined as $\phi=4\eta/3\tau$. The constants $a$ and $b$ are defined as;
$a=\frac{\pi^{2}}{90}\left[16+\frac{21}{2}N_{f}\right]$ (6)
and
$b=(1+1.70N_{f})\frac{0.342}{(1+N_{f}/6)\alpha_{s}^{2}\ln(\alpha_{s}^{-1})}$ (7)
here $N_{f}=3$, is the number of flavors and $\alpha_{s}$ is the strong coupling constant. The first-order viscous correction term, $\phi=4\eta/(3\tau)$, is used to set the initial condition for Eq. (5). By taking the KSS bound of the shear viscosity to entropy density ratio, $\eta/s=1/(4\pi)$, the initial value of the viscous term at $\tau_{0}$ can be expressed as $\phi_{0}=(1/3\pi)(s_{0}/\tau_{0})$. The initial entropy density ($s_{0}$) is estimated using the quasi-particle model (QPM) equation of state at zero baryon chemical potential [32].
II.2 Quarkonia in-medium
Quarkonia traversing through the medium are supposed to be affected by the surrounding partonic environments. The influence of the medium on quarkonia depends on how fast quarkonia is moving in the medium and the relative velocity between the medium and quarkonia. Apart from this, quarkonia, being heavy, do not share the same temperature as the medium. Therefore, to account for the precise impact of the medium on quarkonia, it becomes necessary first to find the effective temperature of the quarkonia within the medium.
II.2.1 Effective Temperature
As mentioned, quarkonium states do not thermalize with the surrounding partonic medium due to their heavy mass and velocities. Therefore, the temperature of the quarkonia in its rest frame can be effectively described through the relativistic Doppler shift (RDS), which arises from the relative motion between the quarkonia and the thermal medium. The resulting effective temperature is quantified in terms of the relative velocity $v_{r}$ between the quarkonium states and the local fluid cell of the medium. The velocities of the medium and the quarkonium states are denoted by $v_{m}$ and $v_{Q(nl)}$, respectively. The latter is related to the quarkonium transverse momentum ($p_{\rm T}$) given as $v_{Q(nl)}=p_{\rm T}/E_{\rm T}$, where $E_{\rm T}=\sqrt{p_{\rm T}^{2}+M_{Q(nl)}^{2}}$ and $M_{Q(nl)}$ is the mass of the quarkonium states. The medium thermal velocity $v_{m}$ is obtained using the Maxwell-Jüttner distribution [34]. The RDS effect leads to an angle-dependent effective temperature experienced by the quarkonium, which is given as [32, 33, 34];
$T_{\text{eff}}(\tau,b,p_{\rm T})=\frac{T(\tau,b)\sqrt{1-|v_{r}|^{2}}}{1-|v_{r}|\cos\theta},$ (8)
where, $\theta$ denotes the angle between $v_{m}$ and $v_{Q(nl)}$. The medium temperature $T(\tau,b)$ is obtained from the solution of Eqs. (4) and (5). To obtain an isotropic (angle-independent) representation of the effective temperature, Eq. (8) is averaged over the solid angle, yielding the mean effective temperature as;
$T_{\text{eff}}(\tau,b,p_{\rm T})=T(\tau,b)\frac{\sqrt{1-|v_{r}|^{2}}}{2|v_{r}|}\ln\left(\frac{1+|v_{r}|}{1-|v_{r}|}\right)$ (9)
The effective temperature of quarkonia encapsulates the modification of the local thermal environment perceived by the moving quarkonium in a relativistically expanding medium.
Figure 1: (Color online) The effective temperature ($T_{\rm eff}$) as a function of $p_{\rm T}$ for $J/\psi$, $\psi$(2S), $\Upsilon$(1S), and $\Upsilon$(2S) states calculated at medium temperature $T$ = 0.400 GeV.
The $T_{\rm eff}$ at $T=400$ MeV as a function of $p_{\rm T}$ for different quarkonium states is characterized in Fig. 1. The initial increase in $T_{\rm eff}$ at $p_{\rm T}\leq M_{Q(nl)}$ arises from the effective blue-shift induced when the quarkonium velocity is much smaller than the medium flow velocity, $v_{Q}\ll v_{m}$. In this regime, the local thermal distribution sampled by the quarkonium is shifted toward higher apparent temperatures due to the relative motion between the probe and the expanding medium. At $v_{Q}\gg v_{m}$, corresponding to the $p_{\rm T}\gg M_{Q(nl)}$, the kinematic red-shift dominates, reducing the apparent slope of the thermal spectrum and consequently decreasing $T_{\rm eff}$ for both charmonium and bottomonium states. This systematic behavior indicates that $T_{\rm eff}$ encodes the interplay between quarkonium kinematics and medium flow, and therefore has a direct impact on the in-medium decay width, and eventually influences the spin alignment of quarkonia as it depends on particle decay width (discussed in Sec. II.3).
II.2.2 Quantum characteristic in presence of rotation
The dynamics of a particle in a rotating medium can be described by a modified Hamiltonian that accounts for the coupling between the spin ($\bf{S}$) of the particle and the angular velocity $\boldsymbol{\omega}$ of the rotating medium. This interaction introduces rotational corrections to the kinetic and potential terms of the system, and the corresponding Hamiltonian can be expressed as [45, 46, 47],
$\mathcal{H}=\frac{1}{2m}\left(\boldsymbol{p}-m\boldsymbol{\omega}\times\mathbf{r}\right)^{2}-\frac{m}{2}(\boldsymbol{\omega}\times\mathbf{r})^{2}-\boldsymbol{\omega}\cdot\mathbf{S}+V(\boldsymbol{r})$ (10)
For a two-body bound system such as quarkonium immersed in a rotating medium, Eq. (10) can be generalized as;
$\mathcal{H}=\sum_{i=1,2}\left[\frac{1}{2m_{i}}(\boldsymbol{p}_{i}-m_{i}\boldsymbol{\omega}_{i}\times\mathbf{r}_{i})^{2}\right$.
[…]
Here, the index $i=1,2$ refers to the heavy quark and its antiquark respectively. Applying the standard reduction procedure for a two-body system to an effective one-body problem in the center of the mass frame, the corresponding canonical momenta and position vectors can be defined as;
$\left\{\begin{array}[]{lcl}\mathbf{P}=\mathbf{p_{1}}+\mathbf{p_{2}},&&\mathbf{p}=m_{\mu}\left(\dfrac{\mathbf{p_{1}}}{m_{1}}-\dfrac{\mathbf{p_{2}}}{m_{2}}\right)\\[8.53581pt] m_{\mu}=\dfrac{m_{1}\;m_{2}}{m_{1}+m_{2}},&\quad&\\[8.53581pt] \mathbf{r_{1}}=\mathbf{R}+\dfrac{m_{2}}{m_{1}+m_{2}}\mathbf{r},&&\mathbf{r_{2}}=\mathbf{R}-\dfrac{m_{1}}{m_{1}+m_{2}}\mathbf{r}\end{array}\right$. (12)
where $m_{\mu}=m_{q}/2$ represents the reduced mass of the heavy quark pair, with the charm and bottom quark masses taken as $m_{c}=1.27$ GeV and $m_{b}=4.18$ GeV, respectively. Considering that both the quark and antiquark experience the same rotation, i.e., $\boldsymbol{\omega}_{1}=\boldsymbol{\omega}_{2}=\boldsymbol{\omega}$, the Hamiltonian simplifies as;
$\mathcal{H}=$ $\;\frac{P^{2}}{2M}+\frac{p^{2}}{2m_{\mu}}-\mathbf{P}\cdot(\boldsymbol{\omega}\times\mathbf{R})-\mathbf{p}\cdot(\boldsymbol{\omega}\times\mathbf{r})$
[…]
Now, focusing on the internal motion in the reduced coordinate system, we obtain;
[…]
Taking rotation along the $z$-axis, i.e., $\boldsymbol{\omega}=\omega\hat{z}$, Eq. (14) can be rewritten as;
[…]
Using the operators $\mathbf{p}=-i\boldsymbol{\nabla}$, $\mathcal{H}=i\frac{\partial}{\partial t}$, $L_{z}=-i\frac{\partial}{\partial\phi}$, and $S_{1z}+S_{2z}=S_{z}$, we can rewrite the Eq. (15) as;
[…]
By expressing the Laplacian operator in spherical coordinates and applying the method of separation of variables, the radial component of the Schrödinger equation can be written as
$\frac{1}{r^{2}}\frac{d}{dr}\Bigl(r^{2}\frac{dR(r)}{dr}\Bigr)+2m_{\mu}\Bigg[(E-V(r))-\frac{l(l+1)}{2m_{\mu}r^{2}}$
$-\omega(L_{z}+S_{z})\Bigg]R(r)=0$ (17)
Now, to include the effect of rotation explicitly in Schrödinger equation, the angular velocity $\omega$ can be expressed in terms of the conserved circulation parameter $C$, defined as,
$C=\oint\boldsymbol{v}\cdot\boldsymbol{dl}$ (18)
Using Stokes’ theorem, which relates the line integral to the surface integral of vorticity, Eq. (18) can be rewritten as;
$C=\oint\boldsymbol{\nabla}\times\boldsymbol{v}\cdot\boldsymbol{dS}=2\omega\pi r^{2}$ (19)
here, $\boldsymbol{\omega}=\frac{1}{2}\boldsymbol{\nabla}\times\boldsymbol{v}$ represents the non-relativistic vorticity. The rotational motion is treated classically and can be related to hydrodynamic vorticity quantities such as kinematic, thermal, temperature, and enthalpy vorticity.
Substituting the expression for $\omega$ in terms of $C$ into Eq. (17) yields the modified radial Schrödinger equation in the rotating frame:
$\frac{1}{r^{2}}\frac{d}{dr}\left(r^{2}\frac{dR(r)}{dr}\right)+2m_{\mu}\left[(E-V(r))-\frac{l(l+1)}{2m_{\mu}r^{2}}\right$.
$\left.-\frac{m_{j}C}{2\pi r^{2}}\right]R(r)=0$ (20)
Here, $m_{j}$ is the magnetic quantum number that arises due to the coupling between $L_{z}$ and $S_{z}$, i.e., $L_{z}+S_{z}=J_{z}$. For a vector meson like quarkonium states $m_{j}=+1,0,-1$.
We have numerically solved the Schrödinger equation by incorporating the real part of the medium modified potential, $V(r,m_{D})$, as given by Eq. (21) and obtained the eigen energies and wave functions for the corresponding quarkonium states. Furthermore, using these variables, we quantified the quarkonia dissociation mechanisms in the medium, which are discussed briefly in the following section.
II.2.3 Collisional Damping
The collisional energy loss of quarkonium in the QGP is commonly referred to as collisional damping, which arises due to the interactions of the bound heavy quark pair with the thermal partons in the medium. This effect leads to a finite in-medium width of the quarkonium state, reflecting its partial decoherence and eventual dissociation. The dissociation rate associated with this mechanism can be evaluated using the imaginary part of the potential, $V(r,m_{D})$, which encodes the medium-induced damping effect. The in-medium singlet complex potential for the heavy quark-antiquark bound state is given as [48, 35]:
$V(r,m_{D})=\frac{\sigma}{m_{D}}(1-e^{-m_{D}\,r})-\alpha_{\rm eff}\left(m_{D}+\frac{e^{-m_{D}\,r}}{r}\right)\\ -i\alpha_{\rm eff}{T_{\rm eff}}\int_{0}^{\infty}\frac{2\,z\,dz}{(1+z^{2})^{2}}\left(1-\frac{\sin(m_{D}\,r\,z)}{m_{D}\,r\,z}\right)$ (21)
In Eq. (21), the first two terms on the right-hand side correspond to the screened string and Coulombic interactions, respectively. The third term represents the imaginary part of the potential, which accounts for the collisional or Landau damping of the low-frequency gauge fields that mediate interactions between quark-antiquark pairs. In Eq. (21), $\sigma$ represents the string tension between the quark-antiquark pair and taken as $\sigma=$ 0.192 GeV2. Further, $m_{D}$ is the Debye screening mass, given as;
$m_{D}=T_{\rm eff}\sqrt{4\pi\alpha_{s}^{T}\left(\frac{N_{c}}{3}+\frac{N_{f}}{6}\right)}$ (22)
here, $N_{c}=3$, $N_{f}=3$ and $\alpha_{s}^{T}$ is the strong coupling constant at the hard scale obtained under the Hard Thermal Loop (HTL) limit. In general, the leading order strong running coupling constant $\alpha_{s}$ is defined as;
$\alpha_{s}(\mu)=\frac{4\pi}{\beta_{0}\ln\left(\frac{\mu^{2}}{\Lambda_{\overline{MS}}^{2}}\right)}$ (23)
where $\beta_{0}=11-\frac{2}{3}N_{f}$, $\Lambda_{\overline{MS}}$ = 0.176 GeV, and $\mu$ is renormalization scale. The $\alpha_{s}^{T}$ is obtained by scaling at the HTL limit ($\mu$ = $2\pi T$), i.e. $\alpha_{s}^{T}=\alpha_{s}(2\pi T)$. As it can be seen from Eq. (23), the $\alpha_{s}^{T}$ depends on medium temperature. At a particular temperature, say T = 400 MeV, $\alpha_{s}^{T}\simeq$ 0.26. Likewise, the coupling constant at the soft scale, $\alpha_{s}^{s}$ and ultra soft scale $\alpha_{s}^{u}$ are determined by evaluating the running coupling at the characteristic scales associated with heavy-quark dynamics; $\alpha_{s}^{s}=\alpha_{s}(m_{q}\alpha_{s}/2)$ and $\alpha_{s}^{u}=\alpha_{s}(m_{q}\alpha_{s}^{2}/2)$. Here, $m_{q}$ is the mass of the charm/bottom quark. Following Eq. (21), effective coupling entering the potential is defined in terms of soft-scale coupling as $\alpha_{\mathrm{eff}}=\frac{4}{3}\alpha_{s}^{s}$. For the chosen set of parameters, we obtain $\alpha_{s}^{s}\simeq 0.33$ for bottom quark, and $\alpha^{s}_{s}\simeq 0.5$ for charm quark, which fixes the numerical value of $\alpha_{\mathrm{eff}}$, respectively.
The collisional damping width, $\Gamma_{\rm damp,nl}$, quantifies the in-medium decay rate of quarkonium due to the imaginary part of the potential. It is computed using first-order perturbation theory as an expectation value of $\mathrm{Im}(V)$ with the radial wave function of the bound state:
$\Gamma_{{\rm damp},nlm}(\tau,p_{\rm T})=\int[g_{nlm}(r)^{\dagger}\left[\rm Im(V)\right]g_{nlm}(r)]dr,$ (24)
here $g_{nlm}(r)$ is the quarkonium singlet wave function. The spin-dependent wave functions for $J/\psi$, $\psi$(2S), $\Upsilon$(1S) and $\Upsilon$(2S) are obtained by solving the Schrödinger equation (Eq. (20)) in the presence of rotation, as described in the preceding section.
II.2.4 Gluonic Dissociation
The gluon-induced dissociation (or gluonic dissociation) of quarkonia in the QGP medium is another dissociation mechanism. This process arises when a thermal gluon interacts inelastically with a color-singlet quarkonium state, leading to a transition to a color-octet state, which subsequently breaks apart. The corresponding dissociation cross-section depends on the gluon energy spectrum and temperature profile of the medium. Following the formalism introduced by Peskin and Bhanot [49, 50] and extended to finite-temperature QCD [51, 52, 32, 33], the gluon-dissociation cross section for a quarkonium state is given by,
$\sigma_{{\rm d},nlm}(E_{g})=\frac{\pi^{2}\alpha_{s}^{u}E_{g}}{N_{c}^{2}}\sqrt{\frac{{m_{q}}}{E_{g}+E_{nlm}}}\\ \;\;\;\;\;\times\left(\frac{l|J_{nlm}^{q,l-1}|^{2}+(l+1)|J_{nlm}^{q,l+1}|^{2}}{2l+1}\right)$ (25)
where, $\alpha_{s}^{u}\simeq 0.76$ for bottom quark, and, $\alpha_{s}^{u}\simeq 0.87$ for charm quark. Here, $E_{g}$ is the incident gluon energy, and $E_{nlm}$ denotes the energy eigenvalue of the quarkonium state. The $g_{nlm}(r)$ and $h_{ql^{\prime}}(r)$ are the color-singlet and color-octet wave functions with orbital angular momenta $l$ and $l^{\prime}$, respectively. The octet wave function $h_{\rm ql^{\prime}}(r)$ has been obtained by solving the Schrödinger equation with the octet potential $V_{8}=\alpha_{\rm eff}/8r$. For a given value of $n$ and $l$, the color-octet wave function and the corresponding energy eigenvalue are identical for all three spin projection states $m$. The value of $q$ is determined by using the conservation of energy, $q=\sqrt{m_{q}(E_{g}+E_{nlm})}$.
The probability density $J_{nlm}^{ql^{\prime}}$ encodes the overlap between the initial bound and final continuum states and is computed using the in medium singlet and octet wave functions,
$J_{nlm}^{ql^{\prime}}=\int_{0}^{\infty}dr\;r\;g^{*}_{nlm}(r)\;h_{ql^{\prime}}(r)$ (26)
As quarkonium moves with a finite velocity relative to the medium, the gluon distribution in the quarkonium rest frame gets modified due to the relativistic Doppler shift of gluon energies. Under a Lorentz transformation, the gluon energy in the quarkonium rest frame is modified as $E_{g}^{\prime}=\gamma E_{g}(1+v_{Q}\cos\theta)$, where $\gamma$ is the Lorentz factor, and $\theta$ is the angle between the quarkonium velocity vector and the incoming gluon momentum. Consequently, the modified gluon distribution function in the quarkonium frame is expressed as,
$f_{g}=\left[\exp\!\left(\frac{E_{g}}{T_{\text{eff}}}\,\gamma(1+v_{Q}\cos\theta)\right)-1\right]^{-1}.$ (27)
This modified distribution plays an essential role in determining the gluonic dissociation rate and, consequently, the momentum-dependent dissociation of quarkonia in the QGP. The thermal averaged gluonic dissociation, $\Gamma_{{\rm gd},nlm}$, can then be obtained by folding the cross section with this modified Bose–Einstein distribution of thermal gluons in the particle rest frame [32, 33], as given by;
$\Gamma_{{\rm gd},nlm}(\tau,p_{\rm T},b)=\frac{g_{d}}{4\pi^{2}}\int_{0}^{\infty}\int_{0}^{\pi}\frac{dp_{g}\,d\theta\,\sin\theta\,p_{g}^{2}\sigma_{{\rm d},nlm}(E_{g})}{e^{\{\frac{\gamma E_{g}}{T_{\rm eff}}(1+v_{Q}\cos\theta)\}}-1}$ (28)
here, $p_{T}$ is the transverse momentum of the quarkonium, and $g_{d}=16$ represents the degeneracy factor for the gluons.
Furthermore, we have obtained the net decay width $\Gamma_{\rm D}$ for quarkonia in-medium via taking the sum over Eq. (24), and (28), given as;
$\Gamma_{{\rm D},nlm}=\Gamma_{{\rm damp},nlm}+\Gamma_{{\rm gd},nlm}$ (29)
The combined effects of collisional damping and gluonic dissociation are subsequently incorporated into the computation of the spin alignment observable $\rho_{00}$, allowing us to explore the role of medium-induced spin–vorticity coupling of quarkonia within a rotating QGP.
II.3 Spin alignment of quarkonia
The dissociation probability of $m^{\rm th}$ spin state in the presence of rotation is given by;
${\mathcal{P}}_{m}=\exp\left[-\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\kern-1.22911pt\tau_{0}$}}^{\text{\normalsize$\scriptstyle\tau_{f}$}}$}}\nolimits\Gamma_{{\rm D},nlm}(\tau,p_{\rm T})\;d\tau\right].$ (30)
Here $m$ takes on the values 0,+1,10,+1,-1, representing the three distinct spin projections corresponding to the states 0, ++1, and 1-1, respectively. The $\tau_{0}$ and $\tau_{f}$ are the proper times corresponding to the initial temperature $T_{0}$ and freeze-out temperature $T_{f}$ for the QGP medium. The dissociation rate, denoted as $\Gamma_{{\rm D},nlm}$ is intricately linked to both the rotational dynamics and the specific spin projection states, $m$.
The spin alignment of vector mesons is measured via the elements of the spin density matrix $\rho_{m,m^{\prime}}$, where $m$ and $m^{\prime}$ denote the spin projections along a chosen quantization axis. For a vector meson with spin 1, the spin density matrix is defined as;
$\rho_{m,m^{\prime}}=\frac{\langle A_{m}A^{*}_{m^{\prime}}\rangle}{\sum\limits^{k=+1}_{k=-1}\langle|A_{k}|^{2}\rangle}$ (31)
where $A_{m}$ denotes the production amplitude for the spin projection $m$, and the denominator ensures proper normalization. The density matrix satisfies the normalization condition
${\rm Tr(\rho)}=\sum\limits^{m=+1}_{m=-1}\rho_{m,m}=1$ (32)
The diagonal elements of the spin density matrix are given by
$\rho_{m,m}=\frac{\langle|A_{m}|^{2}\rangle}{\sum\limits^{k=+1}_{k=-1}\langle|A_{k}|^{2}\rangle}$ (33)
The production amplitude $A_{m}$ for the spin projection $m$, is related to the dissociation probability ${\mathcal{P}}_{m}$ via the relation
${\mathcal{P}}_{m}\equiv\langle|A_{m}|^{2}\rangle$ (34)
The $00^{\rm th}$ element of the $3\times 3$ spin density matrix is defined in terms of the $\Gamma_{{\rm D},nlm}$;
$\rho_{00}=\frac{{\mathcal{P}}_{0}}{{\mathcal{P}}_{+}+{\mathcal{P}}_{0}+{\mathcal{P}}_{-}}$ (35)
III Results and Discussions
Figure 2: The total decay width $\Gamma_{\rm D}$ as a function of transverse momentum ($p_{\rm T}$) for $J/\psi$ (upper left), $\psi$(2S) (upper right), $\Upsilon$(1S) (lower left), and $\Upsilon$(2S) (lower right) in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV. Results are presented for three different temperatures, such as $T$ = 0.400 GeV (red dotted dashed line), $T$ = 0.300 GeV (blue dashed line), and $T$ = 0.175 GeV (green solid line), with the circulation parameter $C=0$ fm.
Quarkonium suppression and spin alignment provide complementary insights into the interaction between heavy quark–antiquark pairs and the quark–gluon plasma. Following that, we evaluated how different dissociation processes modify the spin alignment factor $\rho_{00}$ for the $J/\psi$, $\psi$(2S), $\Upsilon$(1S), and $\Upsilon$(2S) states, as defined in Eq. (35). Here, results are obtained by numerically solving the Schrödinger equation, which provides temperature, vorticity, and spin-projection-dependent eigenvalues and wave functions. The imaginary component of the potential, together with the quarkonium wave function, determines the collisional damping width, whereas the binding energy inferred from the eigenvalue governs the gluon-dissociation width. As reported in the earlier findings [35, 34], the collisional damping mechanism dominates over gluonic dissociation across the thermal evolution of the medium.
Figure 3: The total decay width $\Gamma_{\rm D}$ as a function of transverse momentum ($p_{\rm T}$) for $J/\psi$ (upper), and $\psi$(2S) (lower) in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV for three different spin projections $m_{j}=$ 1 (red dotted dashed line), $m_{j}=$ 1 (blue dashed line), $m_{j}=$ 1 (green solid line) states with the circulation parameter $C=1$ fm. The upper (lower) left and right panels present the $J/\psi$ ($\psi$(2S)) decay width for two different temperatures: $T$ = 0.400 GeV and 0.175 GeV, respectively.
Figure 4: The total decay width $\Gamma_{\rm D}$ as a function of transverse momentum ($p_{\rm T}$) for $\Upsilon$(1S) (upper), and $\Upsilon$(2S) (lower) in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV for three different spin projections $m_{j}=$ 1 (red dotted dashed line), $m_{j}=$ 1 (blue dashed line), $m_{j}=$ 1 (green solid line) states with the circulation parameter $C=1$ fm. The upper (lower) left and right panels present the $\Upsilon$(1S) ($\Upsilon$(2S)) decay width for two different temperatures: $T$ = 0.400 GeV and 0.175 GeV, respectively.
The decay width varies with both the magnitude and the direction of the vorticity field (quantified in terms of the circulation parameter $C$) relative to the quarkonium spin projection. Along with that, it also strongly depends on the transverse momentum ($p_{\rm T}$) and effective temperature ($T_{\rm eff}$) of the particle. It is noteworthy to mention that characteristics of $\Gamma_{\rm D}$ as a function of $p_{\rm T}$ play a crucial role in shaping alignment in the final particle distributions.
[…]
The decay width illustrated in Fig. 2 shows the $p_{\rm T}$ dependence for $J/\psi$ (upper left), $\psi$(2S) (upper right), $\Upsilon$(1S) (lower left), and $\Upsilon$(2S) (lower right) in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV in the absence in-medium rotation ($C=0$ fm) at $T$ = 175, 300, and 400 MeV. The results indicate a strong sensitivity of the dissociation rate to both temperature and momentum. At $T$ = 400 MeV, the $J/\psi$ exhibits an enhanced dissociation rate, which gradually decreases with increasing $p_{\rm T}$. The sudden drop in $\Gamma_{\rm D}$ around $p_{\rm T}\sim$ 10 GeV/c reflects that $T_{\rm eff}\ll T=400$ MeV. Further, at $T$ = 300, and 175 MeV, the effect of $T_{\rm eff}$ significantly diminishes and $\Gamma_{\rm D}$ decreases as a function of $p_{\rm T}$. A similar trend is evident for other quarkonium states, including $\psi$(2S), $\Upsilon$(1S), and $\Upsilon$(2S). In particular, at temperatures of 300 and 175 MeV, $\psi$(2S) exhibits a higher dissociation probability than $J/\psi$ across the entire $p_{\rm T}$ range. This can be attributed to the comparatively lower binding energy of $\psi$(2S) relative to $J/\psi$. Consequently, the reduced survival probability of $\psi$(2S) further supports the sequential melting scenario in the QGP medium. Furthermore, due to the higher mass of $\Upsilon$(1S) and $\Upsilon$(2S) states, which experience relatively stable effective temperatures as they traverse through the medium, as illustrated in Fig. 1. As a result, the dissociation probability for $\Upsilon$(1S) exhibits a modest decreasing trend towards higher $p_{\rm T}$. In contrast, for $\Upsilon$(2S), a slight increase in dissociation probability is observed with rising $p_{\rm T}$. Overall $p_{\rm T}$ dependence of $\Gamma_{\rm D}$ reflects the combined effects of the quarkonium decay width and its effective in-medium temperature. These contrasting behaviors offer important insights into the interplay between quarkonium binding energy and medium effects in heavy-ion collisions.
Further, Fig. 3 and Fig. 4 demonstrate the impact of medium rotation on the $\Gamma_{\rm D}$, as a function of $p_{\rm T}$ for charmonium and bottomonium states, respectively. In the presence of a vortical background, the degeneracy in a quarkonium state is caused by spin-vorticity coupling. This implies that for $m_{j}=\pm 1$ state, the change in net $\Gamma_{\rm D}$ is prominent, which is an artifact of the term ($\sim m_{j}C/r^{2}$) appearing in the effective Hamiltonian. The spin–vorticity coupling alters the eigen energies and radial wave functions in a manner that depends on the sign of $m_{j}$. The spin-vorticity coupling term ($\sim m_{j}C/r^{2}$) is expected to alter the short-distance ($r\rightarrow$ 0 ) behavior of quarkonium wave functions. At a large distance, the behavior of quarkonium wave functions is negligible. For the QCD vacuum potential with $m_{j}=+1$, the additional term due to spin-orbit coupling lowers the effective potential barrier, thereby increasing the spatial extent of the wave function, or making it more peaked at short distances. Consequently, the dissociation probability rises, leading to a noticeably larger total decay width $\Gamma_{\rm D}$ compared to the non-rotating case. In contrast, for $m_{j}=-1$, the rotational term carries the opposite sign, effectively making the potential slightly more confining. This suppresses the sensitivity of the quarkonium wave function at short distances. For the $m_{j}=0$ state, the quarkonium wave function is not affected due to the direct coupling to the rotation term, and naturally follows a trend similar to hydrogen atom wave functions. However, with the medium-modified color-singlet potential for quarkonia, the nature of the quarkonium wave function strongly depends on temperature and effective coupling constant. Further, the dissociation of the quarkonium state in medium is led by the imaginary part of the potential, which induces the Landau damping. Therefore, the net decay width of a quarkonium state is a complex interplay between the temperature-dependent quarkonium wave function and its imaginary part of the complex potential. It is observed that the behavior of the quarkonium wave function at $m_{j}=-1$ and $m_{j}=0$ is almost similar. As a result, the net decay widths for the $m_{j}=0$ and $m_{j}=-1$ states remain closely aligned for both charmonium and bottomonium.
Meanwhile, the splitting between the $m_{j}=+1$ and the other two spin projections for $\Gamma_{\rm D}$ reflects the intrinsic asymmetry introduced by the $m_{j}C$ term in a rotating medium. This rotational effect is more pronounced for charmonium than bottomonium because the latter has a stronger binding, which reduces the relative impact of rotational distortions of the wave function. Nonetheless, for most cases, medium rotation preferentially enhances the dissociation of the $m_{j}=+1$ state in comparison with $m_{j}=0$ and $m_{j}=-1$ states. Furthermore, the change in the $\Gamma_{\rm D}$ pattern with $p_{\rm T}$ in Fig. 3 and Fig. 4 is arises due to the additional contribution of $T_{\rm eff}$ as function of $p_{\rm T}$, which follows the similar explanation as Fig. 2.
Figure 6: (Color online) The spin alignment observable $\rho_{00}$ as a function of transverse momentum ($p_{\rm T}$) for $J/\psi$ (left panel) and $\psi$(2S) (right panel) in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV for various values of the circulation parameter, $C$ in three multiplicity classes.
Figure 5 illustrates $00^{\rm th}$-elements of the spin density matrix, $\rho_{00}$, as a function of charged particle multiplicity ($\langle dN_{\rm ch}/d\eta\rangle$) for $J/\psi$ (upper left), $\psi$(2S) (upper right), $\Upsilon$(1S) (lower left), and $\Upsilon$(2S) (lower right) in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV at mid-rapidity. We have obtained the charged particle multiplicity dependent $\rho_{00}$ for quarkonium state by averaging over $p_{\rm T}$ in the transverse momentum $p_{\rm T}$ range 1 $\leq p_{\rm T}\leq$ 30 GeV/c with the $p_{\rm T}$-distribution function $1/E_{\rm T}^{4}$ [33]. The $p_{\rm T}$-integrated charged particle multiplicity dependent spin alignment is given as,
$\rho_{00}(b)=\frac{\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\kern-1.22911ptp_{\rm Tmin}$}}^{\text{\normalsize$\scriptstyle p_{\rm Tmax}$}}$}}\nolimits dp_{\rm T}\;\rho_{00}(p_{\rm T},b)/(p_{\rm T}^{2}+M_{nl}^{2})^{2}}{\mathop{\text{\large$\int_{\text{\normalsize$\scriptstyle\kern-1.22911ptp_{\rm Tmin}$}}^{\text{\normalsize$\scriptstyle p_{\rm Tmax}$}}$}}\nolimits dp_{\rm T}/(p_{\rm T}^{2}+M_{nl}^{2})^{2}}$ (36)
The results corresponding to observable $\rho_{00}$ are presented for various values of the circulation parameter $C$, which represents the strength of the rotation of the medium. In the absence of the experimental constraint on the $C$, we chose a set of $C$ values spreading from 0.01 fm to 5.0 fm. These chosen values represent the theoretical uncertainty for the prediction of $\rho_{00}$. The dashed horizontal line at $\rho_{00}=1/3$ corresponds to an unpolarized scenario and serves as the baseline for comparison. In the $C=0$ case, corresponding to an irrotational medium, all quarkonium states yield $\rho_{00}=1/3$ independent of charged particle multiplicity, as expected in the absence of spin-vorticity coupling. Finite values of $C$ induce a clear deviation from this baseline, indicating the medium rotation leads to spin-dependent dissociation and consequently generates spin alignment. Now at non-zero $C$, the $\rho_{00}$ for 1S states like $J/\psi$ and $\Upsilon$(1S), remains above $1/3$ over the entire multiplicity range for small and intermediate values of $C$, implying a relatively less dissociation of $m_{j}=0$ spin state. At low multiplicities, where the medium temperature is relatively smaller, the effect of vorticity is more pronounced and therefore, $\rho_{00}$ increases with $\langle dN_{\rm ch}/d\eta\rangle$. As the multiplicity increases, the rise in-medium temperature enhances the relative dissociation corresponding to $m_{j}=\pm 1$ degenerate states. While the effect of spin–vorticity coupling induced due to small values of C, remains marginal and becomes almost ineffective at higher temperatures or higher multiplicities. The combined dynamics of dissociation of $m_{j}=0,\text{and}\pm 1$ degenerate states, along with spin–vorticity coupling respective to small $C$ values, leads to a saturation $\rho_{00}$ at high multiplicities. For larger values of $C$, the enhanced splitting among the spin-dependent decay widths increases the dissociation of the $m_{j}=0$ state, causing $\rho_{00}$ to approach the unpolarized limit for $J/\psi$ and a change in spin alignment orientation for $\Upsilon$(1S) at high multiplicities.
Further, Fig. 5 depicts that quarkonium 2S states like $\psi$(2S) and $\Upsilon$(2S) exhibit $\rho_{00}<1/3$ across the full multiplicity range for all nonzero values of $C$. This behavior reflects the comparable small binding of the $2S$ states, which makes them more susceptible to the thermal effects of the medium rather than rotational effects. As a consequence, the medium temperature dominates over rotational effects except for the substantially large values of C, and leads to enhanced dissociation of the $m_{j}=0$ component relative to the $m_{j}=\pm 1$ states. The $\rho_{00}$ for quarkonium $2S$ states shows a marginal dependence of $C$ over the whole range of multiplicity. Consequently, $\rho_{00}$ decreases monotonically for $\psi$(2S) and $\Upsilon$(2S) with increasing charged particle multiplicity, indicating a transverse spin alignment. Conclusively, Fig. 5 demonstrates that the multiplicity dependence of $\rho_{00}$ arises from the complex dynamics between medium vorticity and temperature-driven dissociation. While rotation-induced spin splitting plays a significant role for tightly bound $1S$ quarkonium states, thermal effects dominate for the loosely bound $2S$ states. This state-dependent behavior highlights the sensitivity of quarkonium spin alignment to both the binding energy of the quarkonia and the vortical structure of the QGP.
Figure 7: (Color online) The spin alignment observable $\rho_{00}$ as a function of transverse momentum ($p_{\rm T}$) for $\Upsilon$(1S) (left panel) and $\Upsilon$(2S) (right panel) in Pb–Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV for various values of the circulation parameter, $C$, in three multiplicity classes.
Figure  6 depicts the $p_{\rm T}$-dependent $\rho_{00}$ for $J/\psi$ (left panels) and $\psi$(2S) (right panels) for three multiplicities classes or centralities in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. For the lowest multiplicity class (upper panels), $\rho_{00}$ exhibits a non-monotonic dependence on $p_{\rm T}$. At low transverse momentum, $\rho_{00}$ remains above the unpolarized limit, exhibiting a reduced dissociation of the $m_{j}=0$ state. With increasing $p_{\rm T}$, $\rho_{00}$ increases and reaches a maximum around $p_{\rm T}\simeq 15$ GeV/$c$, followed by a gradual decrease toward $\rho_{00}=1/3$ at higher $p_{\rm T}$. This behavior directly mirrors the $p_{\rm T}$ dependence of the net decay width $\Gamma_{\rm D}$, which is governed by the interplay between collisional damping, gluonic dissociation, and the momentum-dependent effective temperature. For higher multiplicity classes (middle and lower panels), the $p_{\rm T}$ dependence of $\rho_{00}$ develops a characteristic structure at low transverse momentum. In this region, $\rho_{00}$ initially decreases with increasing $p_{\rm T}$ up to $p_{\rm T}\sim 3$ GeV/$c$, followed by a rising trend at intermediate and high $p_{\rm T}$. This feature is an artifact of the $T_{\rm eff}$. Since $T_{\rm eff}$ enters explicitly into the imaginary part of the in-medium potential, this non-monotonic behavior is directly transmitted to $\Gamma_{\rm D}$ and consequently to $\rho_{00}$.
For a fixed multiplicity class, increasing the circulation parameter $C$ systematically reduces $\rho_{00}$ across the entire $p_{\rm T}$ range without modifying the overall shape of the distribution. This reflects the fact that spin–vorticity coupling alters the relative dissociation probabilities of the spin-projected states through the $m_{j}C$ term in the effective Hamiltonian, while the dominant $p_{\rm T}$ dependence continues to be driven by $T_{\rm eff}$ and $\Gamma_{\rm D}$. The influence of vorticity remains modest for $C\lesssim 1$ fm and becomes more pronounced only at larger values of $C$.
The $\psi$(2S) results shown in the right panels of Fig. 6 demonstrate that for all multiplicity classes, $\rho_{00}^{\psi(2S)}$ remains below the limit $\rho_{00}=1/3$ over the entire $p_{\rm T}$ range, implying a preferential dissociation of the $m_{j}=0$ spin state. The $p_{\rm T}$ dependence of $\rho_{00}$ exhibits a characteristic non-monotonic structure, particularly at higher multiplicities. At low transverse momentum, $\rho_{00}$ decreases with increasing $p_{\rm T}$ up to $p_{\rm T}\sim 10$–15 GeV/$c$, followed by an increase at $p_{\rm T}\geq 15$ GeV/c. This behavior reflects the comparatively weaker binding of the $\psi$(2S) state and sensitivity to the $T_{\rm eff}$, which makes it more susceptible to medium-induced dissociation effects. As a result, temperature-driven dissociation for $\psi$(2S) dominates over rotational effects, leading to a reduced survival probability of the $m_{j}=0$ component relative to the $m_{j}=\pm 1$ states. It consequently shapes the spin alignment of $\psi$(2S) over low to high multiplicities. For a fixed multiplicity class, the dependence of $\rho_{00}$ for $\psi(2S)$} on the circulation parameter $C$ remains weak across the entire $p_{\rm T}$ range. It reflects that spin-vorticity coupling for $\psi$(2S) provides only a subleading correction to the dominant thermal dissociation mechanism. Even at larger values of $C$, the modification of the spin-dependent decay widths is insufficient to overcome the strong temperature sensitivity associated with the small binding energy of the $\psi$(2S) state.
In Fig. 7 spin alignment of $\Upsilon$(1S) and $\Upsilon$(2S) as a function of $p_{\rm T}$ are shown for chosen $C$ values and multiplicity classes in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. At low multiplicity, $\rho_{00}$ for $\Upsilon$(1S) increases with $p_{\rm T}$, consistent with a reduction of the net decay width $\Gamma_{\rm D}$ as the effective temperature $T_{\rm eff}$ decreases at higher transverse momentum. At high multiplicities, $\rho_{00}$ decreases up to $p_{\rm T}\leq 10$ GeV, followed by a gradual increase toward $p_{\rm T}>10$ GeV. It reflects the non-monotonic $p_{\rm T}$ dependence of $T_{\rm eff}$, which enhances dissociation at low transverse momentum and reduces at higher $p_{\rm T}$. The systematic enhancement of $\rho_{00}$ with increasing $C$ for all multiplicity classes indicates that spin–vorticity coupling remains effective for the tightly bound $\Upsilon$(1S) state. In contrast, the $\Upsilon$(2S) state exhibits $\rho_{00}<1/3$ across the whole $p_{\rm T}$-range, and demonstrates the marginal dependence on both $p_{\rm T}$ and $C$. The smaller binding energy of the $\Upsilon$(2S) state enhances its sensitivity to thermal dissociation, which dominates over rotational effects and leads to a preferential dissociation of the $m_{j}=0$ component. Consequently, variations in $T_{\rm eff}$ primarily govern the observed $p_{\rm T}$ dependence for $\Upsilon$(2S), while spin-vorticity coupling introduces only marginal modifications.
Meanwhile, Fig. 7 demonstrates that the $p_{\rm T}$-dependent evolution of spin alignment for bottomonium states governed by the counterintuitive balance between temperature-driven dissociation and rotational effects. While the strongly bound $\Upsilon$(1S) state retains sensitivity to medium vorticity, the more weakly bound $\Upsilon$(2S) state is dominated by thermal dynamics, resulting in a qualitatively different and less vorticity-sensitive spin alignment pattern.
IV Summary
In this work, we have systematically investigated the spin alignment of charmonium ($J/\psi$, $\psi$(2S)) and bottomonium ($\Upsilon$(1S), $\Upsilon$(2S)) states in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV, focusing on dissociation-driven mechanisms in a rotating quark–gluon plasma medium. By incorporating spin–vorticity coupling within an effective Hamiltonian framework and employing a medium-modified complex potential, we evaluated spin-dependent decay widths arising from collisional damping and gluonic dissociation. The space-time evolution of the medium is modeled using second-order relativistic viscous hydrodynamics, and the temperature cooling profile and its impact on quarkonium dynamics are investigated.
Our results demonstrate that medium vorticity induces a spin-dependent modification of the quarkonium decay widths through the $m_{j}C$ term in the effective Hamiltonian, which directly translates into a measurable deviation of the spin density matrix element $\rho_{00}$ from the unpolarized baseline. For tightly bound $1S$ states, $J/\psi$ and $\Upsilon$(1S), $\rho_{00}$ predominantly exceeds $1/3$, indicating a relative stability of the $m_{j}=0$ spin projection and a clear sensitivity to rotational effects, particularly at low to intermediate charged particle multiplicities. In contrast, the excited $2S$ states, $\psi$(2S) and $\Upsilon$(2S), exhibit $\rho_{00}<1/3$ across most of the explored kinematic range, reflecting their enhanced susceptibility to temperature-driven dissociation and a reduced sensitivity to medium vorticity. Further, $p_{\rm T}$-dependence of $\rho_{00}$ reveals a nontrivial structure governed by the momentum-dependent effective temperature $T_{\rm eff}$, which enters explicitly into the imaginary part of the in-medium potential. The resulting non-monotonic behavior of the net decay width $\Gamma_{\rm D}$ with $p_{\rm T}$ is directly imprinted onto the spin alignment observable for both charmonium and bottomonium states. A consistent hierarchy emerges across all observables, wherein the interplay between binding energy, effective temperature, and rotational dynamics determines the final spin alignment pattern.
Conclusively, this study establishes spin-dependent dissociation in a vortical QGP as a viable microscopic mechanism contributing to quarkonium spin alignment. The results provide a unified interpretation of the observed state, momentum, and multiplicity-dependent behavior of $\rho_{00}$, offering a complementary perspective to spin alignment mechanisms based solely on spin transport or hadronization effects. These findings strengthen the role of quarkonium spin observables as sensitive probes of both the thermal and vortical properties of deconfined QCD matter.
Future Outlook
The extension of the present work can further advance the understanding of quarkonium spin dynamics in relativistic heavy-ion collisions.
  • A most general improvement involves embedding the current framework into a 3+1D viscous hydrodynamic, which would allow a more realistic treatment of spatial vorticity fluctuations and local temperature gradients. Such an approach would enable a quantitative assessment of the sensitivity of $\rho_{00}$ to the vorticity.
  • In addition, extending the present analysis to small collision systems, such as high-multiplicity $p$-Pb and ppp-p collisions, would be particularly valuable in light of recent experimental indications of collective behavior. A comparative study across system sizes could disentangle genuine medium-induced effects from initial-state contributions.
  • From an experimental perspective, differential measurements of quarkonium spin alignment in multiple polarization frames and with finer binning in transverse momentum and centrality will be crucial for constraining theoretical models. The present results provide concrete predictions for the state-dependent sensitivity of $\rho_{00}$ to medium rotation, which can be tested in future high-statistics datasets from the LHC and RHIC facilities.
  • Finally, incorporating additional microscopic effects, such as regeneration, feed-down contributions, and possible spin-dependent recombination mechanisms, would further allow a more comprehensive description of quarkonium spin alignment across the evolution of the medium. Together, these developments will help establish quarkonium spin observables as precision tools for probing the vortical structure and microscopic dynamics of the quark–gluon plasma.
Acknowledgement
Bhagyarathi Sahoo acknowledges the financial aid from CSIR, Government of India. The authors gratefully acknowledge the DAE-DST, Government of India, funding under the mega-science project ”Indian Participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021-IITI (E-37123).